Suppose you are playing two unfair coins. The probability of tossing a tail is θ for coin 1, and 2θ for coin 2. You toss each coin for several times, and you get the following results:

(a) What is the probability of tossing a head for coin 1 and for coin 2?

coin no. result

1

head
2 head
1 tail
1 head
2 tail
2 tail

(b) What is the likelihood of the data given \theta?

(c) What is the maximum likelihood estimation for \theta?
Answer #1

(a)

Probability of tossing a head for coin 1 = 1 - heta

Probability of tossing a head for coin 2 = 1 - 2heta

(b)

The likelihood of the data given heta is,

P(head for coin 1) * P(head for coin 2) * P(tail for coin 1) * P(head for coin 1) * P(tail for coin 2) * P(tail for coin 2)

= (1 - heta) * (1 - 2heta) * heta * (1 - heta) * 2heta * 2heta

L(heta) = 4heta^3 (1 - heta)2(1 - 2heta)

(c)

For maximum likelihood estimation of heta

dL(heta)/dheta = 0   

-403(1-9)2(1-20) 0 de dö

1202 (1-0)2(1-20) _ 8θ3(1-6)(1-29) _8θ3(1-0)2-0

3(1-6)(1-29)-29(1-29)-29(1-0) = 0 if heta e 0 and θメ1

Rightarrow 3 + 6 heta^2 - 9 heta - 2 heta + 4 heta^2 - 2 heta + 2 heta^2= 0

1202 130 +3-0

13 ± V132 _ 4 * 12 * 3 2 12

θ= 3/8.1/6

For θ = 318

L( heta) = 4 (3/8)^3 (1-3/8)^2 (1-2*3/8) = 0.02059937

For heta = 1/6

L(0) 4(1/6)3(1-1/6)2(1-2 * 1/6) 0.008573388

Thus, likelihood is maximum for θ = 318

maximum likelihood estimation of heta is 3/8